OpenStudy (anonymous):

CALCULATE NUMBER of molecules in 22.4 litres of water at STP

OpenStudy (ipwnbunnies):

This is a special case that's good to know, I think. At STP, a mole of gas has a volume of 22.4 L. We can show this mathematically with the Ideal Gas Equation: \[PV = nRT\] \[n = \frac{PV}{RT}\] \[n = \frac{(1 atm)(22.4 L)}{(0.08206 L-atm/mol K)(273 K)}\]

OpenStudy (ipwnbunnies):

\[n = 1 mol * 6.02e23 \frac{molecules}{mol} = 6.02e23 molecules\]

OpenStudy (koikkara):

\(\Large~\color{red}{~\star~``Welcome~to~Open~Study~!!"~\star~}\) \(Hope~it~helps~!!\) Don't Forget, \(Appreciate~the~Best~Answer\) with a click on \(\color{purple}{"Best Response"}\) Button Too ! \(\color{darkgreen}{~''Hello"}\) @kriya \(\color{darkgreen}{~''Nice~To~Meet~You~!!"}\) @iPwnBunnies